# Two Hornich-Hlawka-type and Buzano-type inequalities

### v1

@@ Abstract
This paper deals with Hornich-Hlawka-type and Buzano-type inequalities in inner product spaces (IPS). On the one hand, the Hornich-Hlawka inequality, or Hlawka’s inequality, or quadrilateral inequality, is an inequality concerning three vectors in a quadrilateral. Namely, it states that for all $x,y,z$, $\lVert x\rVert+\lVert y\rVert+\lVert z\rVert+\lVert x+y+z\rVert\geq\lVert x+y\rVert+\lVert x+z\rVert+\lVert y+z\rVert$. On the other hand, Buzano inequality, which states $|\langle a,x\rangle\,\langle x,b\rangle|\leq\frac{1}{2}\left[\|a\|\cdot\|b\|+|\langle a,b\rangle|\right]\|x\|^{2}$, is a refinement of the Cauchy-Schwarz inequality, since it is clear that for $a=b$, the above inequality becomes the standard inequality $|\langle a,x\rangle|^{2}\leq\|a\|^{2}\,\|x\|^{2}$. In the first part of our paper, we introduce a stronger version of Hornich-Hlawka inequality, namely: for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In the second part of our paper, we establish a "2-out-of-3" Buzano-type principle.
00footnotetext: 2020 Mathematics Subject Classification. 46C05; 26D15 00footnotetext: Key words and phrases. Hornich-Hlawka inequality; Buzano inequality; inner product spaces; norm inequalities; Cauchy-Schwarz inequality
@@ 1 Introduction
On the other hand, Buzano inequality, which states $|\langle a,x\rangle\,\langle x,b\rangle|\leq\frac{1}{2}\left[\|a\|\cdot\|b\|+|\langle a,b\rangle|\right]\|x\|^{2}$, is a refinement of the Cauchy-Schwarz inequality, since it is clear that for $a=b$, the above inequality becomes the standard inequality $|\langle a,x\rangle|^{2}\leq\|a\|^{2}\,\|x\|^{2}$ (see [1, 11] for other Buzano-type inequalities).
After this introduction, our article is divided into 2 other sections. In section 2, we prove a stronger form of Hornich-Hlawka inequality in inner product spaces, namely: for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In section 3, we prove a ”2-out-of-3” Buzano-type principle. Hereafter, $H$ denotes a (real or complex) inner product space, with scalar product $\langle\cdot,\cdot\rangle$ and norm $\lVert\cdot\rVert$.
@@ 2 A stronger Hornich-Hlawka inequality
    $\bigl(\|x\|+\|y\|+\|z\|+\|x+y+z\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\operatorname{Re}\langle x,y\rangle+\operatorname{Re}\langle y,z\rangle+\operatorname{Re}\langle z,x\rangle\bigr)$
@@
    $\bigl(\|x+y\|+\|y+z\|+\|z+x\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\operatorname{Re}\langle x,y\rangle+\operatorname{Re}\langle y,z\rangle+\operatorname{Re}\langle z,x\rangle\bigr)$
@@
    $L^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\Re{\langle x,y\rangle}+2\Re{\langle x,z\rangle}+2\Re{\langle y,z\rangle}\right)+2\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\lVert z\rVert^{2}\Re{\langle x,y\rangle}+2\lVert z\rVert^{2}\Re{\langle x,z\rangle}+2\lVert z\rVert^{2}\Re{\langle y,z\rangle}$
@@
    $R^{2}$ $=\left(\lVert x\rVert^{2}+\lVert z\rVert^{2}+2\Re{\langle x,z\rangle}\right)\left(\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\Re{\langle y,z\rangle}\right)$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\Re{\langle x,z\rangle}\lVert y\rVert^{2}+2\Re{\langle x,z\rangle}\lVert z\rVert^{2}$
    $\quad\quad+2\Re{\langle y,z\rangle}\lVert x\rVert^{2}+2\Re{\langle y,z\rangle}\lVert z\rVert^{2}+4\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}$
@@
    $\Leftrightarrow\lVert z\rVert^{2}\Re{\langle x,y\rangle}+\lVert z\rVert^{2}\Re{\langle x,z\rangle}+\lVert z\rVert^{2}\Re{\langle y,z\rangle}+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq$
    $\quad\quad\quad\Re{\langle x,z\rangle}\lVert y\rVert^{2}+\Re{\langle x,z\rangle}\lVert z\rVert^{2}+\Re{\langle y,z\rangle}\lVert x\rVert^{2}+\Re{\langle y,z\rangle}\lVert z\rVert^{2}+2\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}$
    $\Leftrightarrow\lVert z\rVert^{2}\Re{\langle x,y\rangle}+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\Re{\langle y,z\rangle}+\lVert y\rVert^{2}\Re{\langle x,z\rangle}+2\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}$
@@
    $\Leftrightarrow\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\Re{\langle y,z\rangle}+\lVert y\rVert^{2}\Re{\langle x,z\rangle}-\lVert z\rVert^{2}\Re{\langle x,y\rangle}+2\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}$
@@
    $a=\lVert x\rVert,\quad b=\lVert y\rVert,\quad c=\lVert z\rVert,\quad p=\Re{\langle x,y\rangle},\quad q=\Re{\langle x,z\rangle},\quad r=\Re{\langle y,z\rangle}.$
@@
    $\textbf{L}^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\Re{\langle x,y\rangle}+2\Re{\langle x,z\rangle}+2\Re{\langle y,z\rangle}\right)$
@@
    $\textbf{R}^{2}$ $=\lVert x\rVert^{4}\Re{\langle y,z\rangle}^{2}+\lVert y\rVert^{4}\Re{\langle x,z\rangle}^{2}+\lVert z\rVert^{4}\Re{\langle x,y\rangle}^{2}+4\Re{\langle x,z\rangle}^{2}\Re{\langle y,z\rangle}^{2}$
    $\quad+2\lVert x\rVert^{2}\lVert y\rVert^{2}\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}-2\lVert x\rVert^{2}\lVert z\rVert^{2}\Re{\langle x,y\rangle}\Re{\langle y,z\rangle}-2\lVert y\rVert^{2}\lVert z\rVert^{2}\Re{\langle x,y\rangle}\Re{\langle x,z\rangle}$
    $\quad+4\lVert x\rVert^{2}\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}^{2}+4\lVert y\rVert^{2}\Re{\langle y,z\rangle}\Re{\langle x,z\rangle}^{2}-4\lVert z\rVert^{2}\Re{\langle x,y\rangle}\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}$
@@
    There exist (complex) numbers $\lambda,\mu$ such that
@@
    There exist (complex) numbers $\lambda,\mu$ such that
@@
    $\|y\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\Re\langle y,z\rangle=0.$
@@
    There exist (complex) numbers $\lambda,\mu$ such that
@@
    $\|x\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\Re\langle x,z\rangle=0.$
@@
$G=\begin{pmatrix}\|x\|^{2}&\Re\langle x,y\rangle&\Re\langle x,z\rangle\\ \Re\langle x,y\rangle&\|y\|^{2}&\Re\langle y,z\rangle\\ \Re\langle x,z\rangle&\Re\langle y,z\rangle&\|z\|^{2}\end{pmatrix}$
@@
3 A ”2-out-of-3” Buzano-type principle
Theorem 3.1 (A 2-out-of-3 Buzano-type principle).
Let $x,y,z\in H$. Assume that
$|\Re{\langle x,z\rangle}|\lVert y\rVert^{2}-|\Re{\langle x,y\rangle}||\Re{\langle y,z\rangle}|\geq 0$
and
$|\Re{\langle x,y\rangle}|\lVert z\rVert^{2}-|\Re{\langle x,z\rangle}||\Re{\langle y,z\rangle}|\geq 0.$
Then necessarily
$|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\langle x,y\rangle||\langle x,z\rangle|\geq 0.$
In particular,
$|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\Re{\langle x,y\rangle}||\Re{\langle x,z\rangle}|\geq 0.$
@@ 3 A ”2-out-of-3” Buzano-type principle
Theorem 3.1 shows that, among the three possible Buzano-type bounds associated with the triple $(x,y,z)$, any two imply the third.
This interesting phenomenon can be interpreted in many ways:
  • It is a single ”weighted Buzano-type inequality”.
  • It is a bridge inequality: if you can control two relationships among three vectors, you automatically control the third. Think of numerical analysis where iterative schemes often produce three consecutive vectors.
  • A lemma for generating new inequalities or for establishing chains of implications between inequalities
  • It bounds cross terms of interacting coupled systems. If two energy inequalities are known (e.g., kinetic and potential parts), the third can often be inferred using results like this.
  • It controls error bounds where only partial inner-product information is available. This is useful when one estimate is hard to prove directly but the other two are easier.
@@
Let $A=\Big\lVert 2\sqrt{|\Re{\langle y,z\rangle}|}x-\frac{|\langle x,y\rangle||\langle x,z\rangle|}{2\sqrt{|\Re{\langle y,z\rangle}|}\Re{\langle x,y\rangle}}y-\frac{|\langle x,y\rangle||\langle x,z\rangle|}{2\sqrt{|\Re{\langle y,z\rangle}|}\Re{\langle x,z\rangle}}z\Big\rVert^{2}\geq 0$. We have
$A$ $=4|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}+\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}}{4|\Re{\langle y,z\rangle}|\Re{\langle x,y\rangle}^{2}}\lVert y\rVert^{2}+\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}}{4|\Re{\langle y,z\rangle}|\Re{\langle x,z\rangle}^{2}}\lVert z\rVert^{2}-2|\langle x,y\rangle||\langle x,z\rangle|-2|\langle x,y\rangle||\langle x,z\rangle|$
$\quad\quad+\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}\Re{\langle y,z\rangle}}{2|\Re{\langle y,z\rangle}|\Re{\langle x,y\rangle}\Re{\langle x,z\rangle}}$
$=4\left[|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\langle x,y\rangle||\langle x,z\rangle|\right]$
$\quad+4\left[\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}\left(|\Re{\langle x,z\rangle}|^{2}\lVert y\rVert^{2}+|\Re{\langle x,y\rangle}|^{2}\lVert z\rVert^{2}+2\Re{\langle x,y\rangle}\Re{\langle x,z\rangle}\Re{\langle y,z\rangle}\right)}{16|\Re{\langle y,z\rangle}||\Re{\langle x,y\rangle}|^{2}|\Re{\langle x,z\rangle}|^{2}}\right]$
which implies
$|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\langle x,y\rangle||\langle x,z\rangle|\geq\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}\left(|\Re{\langle x,z\rangle}|^{2}\lVert y\rVert^{2}+|\Re{\langle x,y\rangle}|^{2}\lVert z\rVert^{2}-2|\Re{\langle x,y\rangle}||\Re{\langle x,z\rangle}||\Re{\langle y,z\rangle}|\right)}{16|\Re{\langle y,z\rangle}||\Re{\langle x,y\rangle}|^{2}|\Re{\langle x,z\rangle}|^{2}}$
$=\frac{|\langle x,y\rangle|^{2}|\langle x,z\rangle|^{2}\left(|\Re{\langle x,z\rangle}|(|\Re{\langle x,z\rangle}|\lVert y\rVert^{2}-|\Re{\langle x,y\rangle}||\Re{\langle y,z\rangle}|)+|\Re{\langle x,y\rangle}|(|\Re{\langle x,y\rangle}|\lVert z\rVert^{2}-|\Re{\langle x,z\rangle}||\Re{\langle y,z\rangle}|)\right)}{16|\Re{\langle y,z\rangle}||\Re{\langle x,y\rangle}|^{2}|\Re{\langle x,z\rangle}|^{2}}$
Therefore, if
$|\Re{\langle x,z\rangle}|\lVert y\rVert^{2}-|\Re{\langle x,y\rangle}||\Re{\langle y,z\rangle}|\geq 0$
and
$|\Re{\langle x,y\rangle}|\lVert z\rVert^{2}-|\Re{\langle x,z\rangle}||\Re{\langle y,z\rangle}|\geq 0$
then
$|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\langle x,y\rangle||\langle x,z\rangle|\geq 0$
In particular,
$|\Re{\langle y,z\rangle}|\lVert x\rVert^{2}-|\Re{\langle x,y\rangle}||\Re{\langle x,z\rangle}|\geq 0$
since $|\langle x,y\rangle|\geq|\Re{\langle x,y\rangle}|$ and $|\langle x,z\rangle|\geq|\Re{\langle x,z\rangle}|$. ∎

### v2

# Two Hornich-Hlawka-type and Gram matrix-based inequalities

@@ Abstract
This paper deals with Hornich-Hlawka-type and Gram matrix-based inequalities in inner product spaces (IPS). On the one hand, the Hornich-Hlawka inequality, or Hlawka’s inequality, or quadrilateral inequality, is an inequality concerning three vectors in a quadrilateral. Namely, it states that for all $x,y,z$, $\lVert x\rVert+\lVert y\rVert+\lVert z\rVert+\lVert x+y+z\rVert\geq\lVert x+y\rVert+\lVert x+z\rVert+\lVert y+z\rVert$. On the other hand, inequalities based on Gram matrices are very important because they generate generically all inequalities implying three vectors in an IPS. In the first part of our paper, we introduce a stronger version of Hornich-Hlawka inequality, namely, for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In the second part of our paper, we establish a sharp inequality based on Gram matrices that implies, in particular, the standard Cauchy-Schwarz inequality.
00footnotetext: 2020 Mathematics Subject Classification. 46C05; 26D15 00footnotetext: Key words and phrases. Hornich-Hlawka inequality; Gram matrix; Cauchy-Schwarz inequality; inner product spaces; norm inequalities
@@ 1 Introduction
On the other hand, important inequalities can be established from the nonnegativity of Gram matrices. Actually, all inequalities implying three vectors in an IPS can be generically derived from the nonnegativity of the Gram matrices of these vectors.
After this introduction, our article is divided into 2 other sections. In section 2, we prove a stronger form of Hornich-Hlawka inequality in inner product spaces, namely: for all vectors $x,y,z$, $\|x\|\|y\|+\|z\|\|x+y+z\|\geq\|x+z\|\|y+z\|$. In section 3, we prove a sharp inequality based on Gram matrices that strengthens, in particular, the Cauchy-Schwarz inequality. Hereafter, $H$ denotes a real inner product space, with scalar product $\langle\cdot,\cdot\rangle$ and norm $\lVert\cdot\rVert$.
@@ 2 A stronger Hornich-Hlawka inequality
    $\bigl(\|x\|+\|y\|+\|z\|+\|x+y+z\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\langle x,y\rangle+\langle y,z\rangle+\langle z,x\rangle\bigr)$
@@
    $\bigl(\|x+y\|+\|y+z\|+\|z+x\|\bigr)^{2}=2\bigl(\|x\|^{2}+\|y\|^{2}+\|z\|^{2}\bigr)+2\bigl(\langle x,y\rangle+\langle y,z\rangle+\langle z,x\rangle\bigr)$
@@
    $L^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle x,y\rangle+2\langle x,z\rangle+2\langle y,z\rangle\right)+2\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\lVert z\rVert^{2}\langle x,y\rangle+2\lVert z\rVert^{2}\langle x,z\rangle+2\lVert z\rVert^{2}\langle y,z\rangle$
@@
    $R^{2}$ $=\left(\lVert x\rVert^{2}+\lVert z\rVert^{2}+2\langle x,z\rangle\right)\left(\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle y,z\rangle\right)$
    $=\lVert x\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{2}\lVert x\rVert^{2}+\lVert z\rVert^{2}\lVert y\rVert^{2}+\lVert z\rVert^{4}+2\langle x,z\rangle\lVert y\rVert^{2}+2\langle x,z\rangle\lVert z\rVert^{2}$
    $\quad\quad+2\langle y,z\rangle\lVert x\rVert^{2}+2\langle y,z\rangle\lVert z\rVert^{2}+4\langle x,z\rangle\langle y,z\rangle$
@@
    $\Leftrightarrow\lVert z\rVert^{2}\langle x,y\rangle+\lVert z\rVert^{2}\langle x,z\rangle+\lVert z\rVert^{2}\langle y,z\rangle+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq$
    $\quad\quad\quad\langle x,z\rangle\lVert y\rVert^{2}+\langle x,z\rangle\lVert z\rVert^{2}+\langle y,z\rangle\lVert x\rVert^{2}+\langle y,z\rangle\lVert z\rVert^{2}+2\langle x,z\rangle\langle y,z\rangle$
    $\Leftrightarrow\lVert z\rVert^{2}\langle x,y\rangle+\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\langle y,z\rangle+\lVert y\rVert^{2}\langle x,z\rangle+2\langle x,z\rangle\langle y,z\rangle$
@@
    $\Leftrightarrow\lVert x\rVert\lVert y\rVert\lVert z\rVert\lVert x+y+z\rVert\geq\lVert x\rVert^{2}\langle y,z\rangle+\lVert y\rVert^{2}\langle x,z\rangle-\lVert z\rVert^{2}\langle x,y\rangle+2\langle x,z\rangle\langle y,z\rangle$
@@
    $a=\lVert x\rVert,\quad b=\lVert y\rVert,\quad c=\lVert z\rVert,\quad p=\langle x,y\rangle,\quad q=\langle x,z\rangle,\quad r=\langle y,z\rangle.$
@@
    $\textbf{L}^{2}$ $=\lVert x\rVert^{2}\lVert y\rVert^{2}\lVert z\rVert^{2}\left(\lVert x\rVert^{2}+\lVert y\rVert^{2}+\lVert z\rVert^{2}+2\langle x,y\rangle+2\langle x,z\rangle+2\langle y,z\rangle\right)$
@@
    $\textbf{R}^{2}$ $=\lVert x\rVert^{4}\langle y,z\rangle^{2}+\lVert y\rVert^{4}\langle x,z\rangle^{2}+\lVert z\rVert^{4}\langle x,y\rangle^{2}+4\langle x,z\rangle^{2}\langle y,z\rangle^{2}$
    $\quad+2\lVert x\rVert^{2}\lVert y\rVert^{2}\langle x,z\rangle\langle y,z\rangle-2\lVert x\rVert^{2}\lVert z\rVert^{2}\langle x,y\rangle\langle y,z\rangle-2\lVert y\rVert^{2}\lVert z\rVert^{2}\langle x,y\rangle\langle x,z\rangle$
    $\quad+4\lVert x\rVert^{2}\langle x,z\rangle\langle y,z\rangle^{2}+4\lVert y\rVert^{2}\langle y,z\rangle\langle x,z\rangle^{2}-4\lVert z\rVert^{2}\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle$
@@
    There exist real numbers $\lambda,\mu$ such that
@@
    There exist real numbers $\lambda,\mu$ such that
@@
    $\|y\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\langle y,z\rangle=0.$
@@
    There exist real numbers $\lambda,\mu$ such that
@@
    $\|x\|^{2}\,\lambda(\lambda+1)+\|z\|^{2}\,\mu(\mu+1)+2\lambda(\mu+1)\langle x,z\rangle=0.$
@@
$G=\begin{pmatrix}\|x\|^{2}&\langle x,y\rangle&\langle x,z\rangle\\ \langle x,y\rangle&\|y\|^{2}&\langle y,z\rangle\\ \langle x,z\rangle&\langle y,z\rangle&\|z\|^{2}\end{pmatrix}$
@@
3 An inequality based on Gram matrices
Theorem 3.1.
We have, for all vectors $x,y,z\in H$ and real numbers $\alpha,\beta,\gamma$:
$\alpha^{2}\|x\|^{2}\langle y,z\rangle^{2}+\beta^{2}\|y\|^{2}\langle x,z\rangle^{2}+\gamma^{2}\|z\|^{2}\langle x,y\rangle^{2}+2(\alpha\beta+\alpha\gamma+\beta\gamma)\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0.$
Proof.
Given $x,y,z$ in $H$, the Gram matrix associated to $(x,y,z)$ is
$G=\begin{pmatrix}\langle x,x\rangle&\langle x,y\rangle&\langle x,z\rangle\\ \langle y,x\rangle&\langle y,y\rangle&\langle y,z\rangle\\ \langle z,x\rangle&\langle z,y\rangle&\langle z,z\rangle\end{pmatrix}$.
It is positive semidefinite, so for any real numbers $\alpha,\beta,\gamma$, we have
$Q(\alpha,\beta,\gamma)=(\alpha,\beta,\gamma)\,G\,\begin{pmatrix}\alpha\\ \beta\\ \gamma\end{pmatrix}\geq 0.$
Expanding, we have
$Q(\alpha,\beta,\gamma)=\alpha^{2}\|x\|^{2}+\beta^{2}\|y\|^{2}+\gamma^{2}\|z\|^{2}+2\alpha\beta\langle x,y\rangle+2\alpha\gamma\langle x,z\rangle+2\beta\gamma\langle y,z\rangle\geq 0.$
Replace
$\alpha\mapsto\alpha\langle y,z\rangle,$
$\beta\mapsto\beta\langle x,z\rangle,$
$\gamma\mapsto\gamma\langle x,y\rangle.$
Then
$R(\alpha,\beta,\gamma)$ $=Q(\alpha\langle y,z\rangle,\beta\langle x,z\rangle,\gamma\langle x,y\rangle)$
$=\alpha^{2}\|x\|^{2}\langle y,z\rangle^{2}+\beta^{2}\|y\|^{2}\langle x,z\rangle^{2}+\gamma^{2}\|z\|^{2}\langle x,y\rangle^{2}$
$\quad+2(\alpha\beta+\alpha\gamma+\beta\gamma)\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0.$
Corollary 3.2.
  • If $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle>0$, we have
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq 3\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
  • If $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle<0$, we have
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq-6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
These inequalities are sharp.
@@ 3 An inequality based on Gram matrices
  • This corollary can be understood as a reverse inequality of the Gram determinant inequality
    $\det G\geq 0$
    which translates to
    $\|x\|^{2}\|y\|^{2}\|z\|^{2}+2\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}.$
  • The first point of the corollary implies the Cauchy-Schwarz inequality, and therefore all other inequalities deduced from it. One can show this by taking $z=y$.
@@
  • If
    $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle\geq 0$
    Then R(1,1,2)>0R(1,1,-2)>0 implies
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+4\|z\|^{2}\langle x,y\rangle^{2}\geq 6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
    Summing up symmetrically and dividing by 6, we get
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq 3\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
  • If
    $\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle<0$
    Then $R(1,1,1)\geq 0$ implies
    $\|x\|^{2}\langle y,z\rangle^{2}+\|y\|^{2}\langle x,z\rangle^{2}+\|z\|^{2}\langle x,y\rangle^{2}\geq-6\langle x,y\rangle\langle x,z\rangle\langle y,z\rangle.$
Sharpness. For $D>0$, take $x=y=z$ nonzero. It yields LHS = RHS.
For $D<0$, take $x=(1,0,0)$, $y=\left(\frac{1}{2},\frac{\sqrt{3}}{2},0\right)$, $z=\left(\frac{1}{2},-\frac{\sqrt{3}}{2},0\right)$.
We have
$\langle x,y\rangle=1\cdot\frac{1}{2}+0\cdot\frac{\sqrt{3}}{2}=\frac{1}{2}$
$\langle x,z\rangle=1\cdot\frac{1}{2}+0\cdot\left(-\frac{\sqrt{3}}{2}\right)=\frac{1}{2}$
$\langle y,z\rangle=\frac{1}{2}\cdot\frac{1}{2}+\frac{\sqrt{3}}{2}\cdot\left(-\frac{\sqrt{3}}{2}\right)=\frac{1}{4}-\frac{3}{4}=-\frac{1}{2}$
So
$RHS=-6\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)\left(-\frac{1}{2}\right)=\frac{6}{8}=\frac{3}{4}.$
Moreover
$\|x\|^{2}\langle y,z\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
$\|y\|^{2}\langle x,z\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
$\|z\|^{2}\langle x,y\rangle^{2}=1\cdot\frac{1}{4}=\frac{1}{4}$
So
$LHS=\frac{1}{4}+\frac{1}{4}+\frac{1}{4}=\frac{3}{4}.$
Equality holds. ∎